# No Query, No Access, Yet Dangerous: Victim Data-based Adversarial Attack

### v1

@@ 1 Introduction
Text classification, a key task in natural language processing, assigns text to categories (Fields et al., 2024). Neural networks, especially large language models (LLMs), have greatly improved performance (Sun et al., 2023). However, adversarial attacks can subtly alter text, causing drastic output changes and undermining system security, with potential financial and legal consequences (Wang et al., 2023; Han et al., 2024; Fursov et al., 2021).
@@ 4.1 Cluster-Guided Shadow Dataset Synthesis
$f_{\text{s}}(\mathbf{x})\neq f_{\text{s}}(\tilde{\mathbf{x}}),\text{and}\quad f_{\text{v}}(\mathbf{x})\neq f_{\text{v}}(\tilde{\mathbf{x}}),$ (3)
@@ 4.3 Diverse Adversarial Example Generation
Figure 5: The overview of Diverse Adversarial Example Generation. The left figure depicts the generation of six candidate adversarial samples using two attack methods ($M1$, $M2$) and three substitute models ($f_{1}^{\text{s}}$, $f_{2}^{\text{s}}$,$f_{3}^{\text{s}}$). According to Equation 9, $\tilde{\mathbf{x}}_{i}^{11}$ and $\tilde{\mathbf{x}}_{i}^{21}$ are eliminated for violating the constraint:$p_{\hat{y}}^{f_{j}^{\text{s}}}(\mathbf{x}_{i}^{lj})-p_{\hat{y}}^{f_{j}^{\text{s}}}(\tilde{\mathbf{x}}_{i}^{lj})\geq\epsilon^{{}^{\prime}}$ which suggests that $\mathbf{x}_{i}$ is close to the decision boundary of $f_{1}^{s}$, rendering it an unsuitable substitute model. Similarly, $\tilde{\mathbf{x}}_{i}^{12}$ and $\tilde{\mathbf{x}}_{i}^{22}$ are excluded due to insufficient similarity, failing to satisfy: $\frac{f_{\text{e}}(\mathbf{x})\cdot f_{\text{e}}(\tilde{\mathbf{x}}_{i}^{lj})}{\|f_{\text{e}}(\mathbf{x})\|\cdot\|f_{\text{e}}(\tilde{\mathbf{x}}_{i}^{lj})\|}\geq\epsilon^{*}$ The remaining candidates, $\tilde{\mathbf{x}}_{i}^{13}$ and $\tilde{\mathbf{x}}_{i}^{23}$, are assessed according to their attack performance scores $s_{ilj}$. The sample with the highest score, $\tilde{\mathbf{x}}_{i}^{13}$, is chosen as the final adversarial example for $\mathbf{x}_{i}$.
@@ 5.3 Ablation Study
To explore factors influencing VDBA attack performance, including cluster numbers, training iterations of the hierarchical substitute model, attack methods, and clustering and embedding techniques, we conduct an ablation study. We evaluated the impact of cluster numbers by testing 2, 3, and 4 clusters. Training iterations were assessed by training substitute models for 1, 2, 3, and 4 iterations. The effect of the number of attack methods was tested with configurations of 1 (FD), 3 (FD, Hotflip, PSO), 5 (BAE, FD, Hotflip, PSO, TextBugger), and 7 (BAE, FD, Hotflip, PSO, TextBugger, LEAP, HQA) methods. Clustering methods were supplemented with K-means (Lloyd, 1982), Spectral Clustering (Von Luxburg, 2007), and BIRCH (Zhang et al., 1996). For embedding, we incorporated CLIP (Radford et al., 2021) and one-hot encoding (Rodríguez et al., 2018), in addition to the T5 pre-trained model. CLIP, trained on image-text pairs, uses a dataset distinct from typical text classification datasets, reducing overlap with the victim model’s data. One-hot encoding, a non-pretrained approach, relies solely on victim texts, avoiding external data interference.
@@ 3. Probability of Maximum Exceeding the Maximum of a Subset
$\lim_{w\to\infty}P^{\text{sim}}_{(w)}\;=\;1\quad\text{and}\quad\mu_{w}\to\infty\;\text{(if }p_{i}^{\text{sim}}\text{ sums to infinity)}.$) .

### v2

@@ 1 Introduction
Text classification, a key task in natural language processing, assigns text to categories (Fields et al., 2024). Neural networks, especially large language models (LLMs), have greatly improved performance (Sun et al., 2023). However, adversarial attacks can subtly alter text, causing drastic output changes and undermining system security, with potential financial and legal consequences (Wang et al., 2023; Han et al., 2024; Fursov et al., 2021; Wei et al., 2018; Liang et al., 2020, 2022c, 2022a; Ying et al., 2024; Jing et al., 2025; Liang et al., 2023, 2024, 2022b; Chen et al., 2024).
@@ 4.1 Cluster-Guided Shadow Dataset Synthesis
[…]
@@ 4.3 Diverse Adversarial Example Generation
[…]
@@ 5.3 Ablation Study
To explore factors influencing VDBA attack performance, including cluster numbers, training iterations of the hierarchical substitute model, attack methods, and clustering and embedding techniques, we conduct an ablation study. We evaluated the impact of cluster numbers by testing 2, 3, and 4 clusters. Training iterations were assessed by training substitute models for 1, 2, 3, and 4 iterations. The effect of the number of attack methods was tested with configurations of 1 (FD), 3 (FD, Hotflip, PSO), 5 (BAE, FD, Hotflip, PSO, TextBugger), and 7 (BAE, FD, Hotflip, PSO, TextBugger, LEAP, HQA) methods. Clustering methods were supplemented with K-means (Lloyd, 1982), Spectral Clustering (Von Luxburg, 2007), and BIRCH (Zhang et al., 1996). For embedding, we incorporated CLIP (Radford et al., 2021) and one-hot encoding (Rodr´ıguez et al., 2018), in addition to the T5 pre-trained model. CLIP, trained on image-text pairs, uses a dataset distinct from typical text classification datasets, reducing overlap with the victim model’s data. One-hot encoding, a non-pretrained approach, relies solely on victim texts, avoiding external data interference.
@@ 3. Probability of Maximum Exceeding the Maximum of a Subset
$\lim_{w\to\infty}P^{\text{sim}}_{(w)}\;=\;1\quad\text{and}\quad\mu_{w}\to\infty\;\text{(if }p_{i}^{\text{sim}}\text{ sums to infinity)}.$